# How many qubits does a machine learning problem require?

### v1

@@ 4 Techniques: Bit Encoding Schemes
Bit-bit encoding maps numeric features and class labels into bitstrings for quantum classification. A necessary aspect of this process is compressing the original dataset. In this section, we empirically investigate how many qubits are required to fully cover (represent) bit-bit encoded machine learning datasets. We consider training and testing sets separately when calculating coverage. We calculate the training set coverage as 1training collision incidence1-$\text${training collision incidence}, where the training collision incidence is the proportion of training samples with different class labels which collide (i.e. have the same encoding). A training set is fully covered when there are no collisions between samples with different class labels (i.e. samples with the same encoding have the same class label). In this case, since there are no colliding samples with different classes, a perfect model would theoretically be able to achieve $100\%$ training accuracy. Therefore, we refer to 1training collision incidence1-$\text${training collision incidence} as the theoretical training accuracy. We define the testing set coverage as 1testing overlap incidence1-$\text${testing overlap incidence}, where the testing overlap incidence is the proportion of testing samples which overlap with the training data but do not have the same class label as the majority class label of the overlapping training data. A testing set is fully covered when all of the samples either 1.) are not overlapping with the training data or 2.) have the same class label as the majority class label of the overlapping training data. We refer to 1testing overlap incidence1-$\text${testing overlap incidence} as the theoretical testing accuracy. The minimum number of qubits at which the training and testing sets are covered (i.e. $100\%$ theoretical accuracy) is $Q_{\text{train}}$ and $Q_{\text{test}}$, respectively. The number of qubits required to cover the class labels is $Q_{y}=\lceil\log_{2}{(\text{number of classes})}\rceil$. The minimum number of qubits to cover an entire dataset is $Q_{\text{dataset}}=\max{(Q_{\text{train}},Q_{\text{test}})}+Q_{y}$. While in the bit-bit encoding algorithm $D$, the number of components after the dimensionality reduction scheme acts on the original features, is a modifiable parameter, in all experiments we set $D$ equal to the original number of features.
4.1 Simple PCA
@@
Figure 1: The train and test theoretical accuracy (left), test-train overlap (middle), and correctly classified test-train overlap (right) over the number of allocated qubits for the MNIST $784$ OpenML dataset with PCA dimensionality reduction. Each line shows the mean of the corresponding metric with $95\%$ confidence interval error bands. The gray dotted line is the mean $Q_{\text{dataset}}$ (minus the class qubits) when the theoretical accuracy threshold is $0.99$, instead of the default $1.0$.
@@
We start with a simple example of calculating $Q_{\text{dataset}}$ for the MNIST digit-recognition dataset. We use the MNIST $784$ dataset from OpenML [31], which converts the 2828x2828 images into $784$ features, where the features are the grayscale intensity of each pixel. The MNIST dataset contains $70,000$ samples and $10$ classes, one class for each single-digit number. Figure 1, left, shows the mean theoretical accuracy across $10$ replicates of training/testing splits. $Q_{\text{dataset}}$ is the number of class qubits plus the number at which both the training and testing theoretical accuracy reach $1$. The maximum $Q_{\text{dataset}}$ for MNIST $784$ is $63$, where the x-axis of the figure is cut-off, while the minimum $Q_{\text{dataset}}$ is $42$ and the mean is $50$. The gray dotted line shows the mean $Q_{\text{dataset}}$ ($27$) minus the number of class qubits ($4$) if the theoretical accuracy threshold was $0.99$ instead of $1.0$. The difference between $Q_{\text{dataset}}$ between thresholds is explained by the theoretical training accuracy slowly increasing from $0.99$ to $1.0$ over the number of qubits. We postulate that the difference between the $Q_{\text{dataset}}$ at the threshold of $0.99$ versus $1.0$ reveals a trade-off between achieving the maximum theoretical training accuracy and obtaining a generalizable testing set; as the number of qubits increases, the overlap between the training and test sets decreases (Figure 1, middle). Furthermore, as seen in Figure 1, right, the dip in the testing accuracy between $10$ and $20$ qubits is caused by shifts in how the unique bit strings overlap across classes as the encoding changes with the number of qubits allocated. Between $20$ and $50$ qubits, test-train overlap is near but not quite 0, which is why there is stochasticity in correctly classified test-train overlap in that range.
@@
Next, we calculate $Q_{\text{dataset}}$ for a variety of benchmark classification datasets from OpenML [31]. We use all OpenML datasets that fulfill the following criteria: less than $1,000$ features which are all numeric, less than $10,000$ samples, more samples than features, no missing values, single-label classification with less than $10$ classes, and no collisions when not encoded. After filtering, we have $29$ datasets. Supplementary Figure S1 visualizes the number of samples, number of features, and number of class labels of each dataset. We randomly split each dataset into $80\%$ training and $20\%$ testing, repeated across $10$ replicates. We obtain $Q_{\text{dataset}}$ of each dataset by calculating the training collision and testing overlap incidence when continuously increasing the number of qubits allocated to the dataset from $1$ until the training and testing collision are both 0. We find the mean $Q_{\text{dataset}}$ for the OpenML datasets to be $20.49$ qubits, $95\%$ CI $[19.97,21.00]$. Supplementary Figure 2 shows the theoretical accuracy for the training and testing sets. We see each dataset converges to its $Q_{\text{dataset}}$ at a different rate and that there is minimal stochasticity in the theoretical accuracy across replicates.
4.2 Alternatives to PCA
@@
We now compare $Q_{\text{dataset}}$ across different dimensionality reduction schemes. We include a null comparison where we don’t transform the data with a dimensionality reduction scheme (NONE). We compare three linear schemes: PCA, Independent Component Analysis (ICA), and truncated singular vector decomposition / latent semantic analysis (LSA). We also compare three non-linear schemes: non-negative matrix factorization (NMF), radial basis function kernel PCA (KPCA), and uniform manifold approximation and projection (UMAP). All schemes, except UMAP [32], are implemented with Scikit-Learn [33]. Some dimensionality reduction schemes require special processing of the data- for NMF and KPCA, the training and testing sets are min-max normalized before reduction, and for ICA, the number of components is set to the minimum between the number of features and the rank of the training data.
@@
Figure 2: Violin plots of $Q_{\text{dataset}}$ for OpenML datasets across dimensionality reduction schemes. The y-axis is cut-off at $100$ for ease of visualization. Each violin depicts the distribution of $Q_{\text{dataset}}$ for a given scheme by showing its probability density across $Q_{\text{dataset}}$ values.
Figure 2 shows the distributions of $Q_{\text{dataset}}$ across dimensionality reduction schemes. Overall, ICA has the lowest rounded mean $Q_{\text{dataset}}$ at $19$, followed by PCA ($20$), LSA ($21$), KPCA ($21$), NMF ($26$), NONE ($49$), and UMAP ($70$). ICA, PCA, LSA, and KPCA aim to reduce statistical redundancy among features. Specifically, PCA, LSA, and KPCA transform the original feature space into orthogonal components, which are uncorrelated, while ICA produces statistically independent components. If components are statistically dependent, information about the class label $y$ may be redundantly captured across multiple components. To see this, let $X=[x_{1},\dots,x_{D}]$ denote the $D$ components after dimensionality reduction. The true information the components provide about the class is given by the joint mutual information $I(X;y)$ [34]. However, the importance score of a feature $i$ is given by $I(X_{i};y)$, the mutual information between feature $i$ and $y$. Therefore, if the components are statistically dependent, then the sum $\sum_{i=1}^{D}{I(X_{i};y)}$ over-counts overlapping information. For example, if $X_{i}$ and $X_{j}$ are statistical dependent ($I(X_{i};X_{j})>0$), then some portion of the information they provide about $y$ could be redundant. In general, the over-count among all components can be written as: $\sum_{i=1}^{D}{I(X_{i};y)}-I(X;y)$, the overestimation that arises from feature dependence. Since the allocation of bits is based on $\sum_{i}I(X_{i};y)$, bit-bit encoding with statistically dependent components distributes too many bits to redundant features and too few to unique, complementary features. In contrast, if the components are independent, then $\sum_{i=1}^{D}{I(X_{i};y)}=I(X;y)$, so there is no redundancy. Thus, dimensionality reduction schemes that produce independent or de-correlated components can produce smaller $Q_{\text{dataset}}$ if the allocation of bits reflects the true joint information.
@@ 4.3 Results on Resource Estimation
Table 1: $Q_{\text{dataset}}$ statistics when the threshold is $0.99$ theoretical accuracy vs $1.0$ for the 55 most-run OpenML datasets that don’t have symbolic features, rounded to the nearest qubit. Each dataset has $10$ replicates with different train/test splits. $Q_{\text{dataset}}$ is calculated with PCA dimensionality reduction.
Dataset Mean $Q_{\text{dataset}}$ $(0.99)$ Mean $Q_{\text{dataset}}(1.0)$
steel-plates-fault 15 20
qsar-biodeg 24 110
wdbc 14 18
phoneme 21 28
diabetes 18 21
@@
The universal approximation property of bit-bit encoding allows us to state that a dataset can be completely modeled with $Q_{\text{dataset}}$ qubits. Hence, we can then say that if $Q_{\text{dataset}}$ is within classical simulation capabilities, $\lessapprox 50$ qubits [21], the corresponding dataset is unlikely to benefit from quantum advantage. Datasets that require $\gtrapprox 50$ qubits are likely promising targets for quantum machine learning within this framework. In Section 4.2, we saw that classical machine learning datasets encoded to minimize statistical dependency between features had $Q_{\text{dataset}}$ of around $20$, on average. These classical machine learning datasets had less than $1,000$ features, less than $10,000$ samples, and less than $10$ classes (Supplementary Figure 1). This result suggests that classic, medium-sized machine learning datasets are unlikely to benefit from quantum advantage. Furthermore, in Section 4.1 we saw a dramatic difference in $Q_{\text{dataset}}$ depending on if the theoretical accuracy threshold at which to record $Q_{\text{train}}$ and $Q_{\text{test}}$ was $0.99$ or $1.0$. In Table 1 we further explore this difference in the top-five most-run OpenML datasets that fulfilled the criteria outlined in Section 4.1. We see, again, a dramatic difference in $Q_{\text{dataset}}$ when the threshold is $0.99$ vs $1.0$, suggesting a trade-off between complete modeling of the training set and generalization on the testing set.
4.4 Schemes for Very Large Datasets
@@
Figure 3: The table shows how the rounded mean $Q_{\text{dataset}}$ of subsamples of the Tahoe dataset scales with the number of features and number of samples.
@@
The default bit-bit encoding algorithm requires the entire dataset to be read into memory, which becomes impractical for large datasets. Furthermore, dimensionality reduction algorithms may not scale well as the dataset size increases, leading to intractable time and memory overhead. For example, PCA dimensionality reduction scales polynomially with dataset size. Learning problems on large datasets are more likely to benefit from quantum advantage, since large datasets should have more bits of information that can only be compressed without collision and overlap with a significant number of qubits. Therefore, we extend the bit-bit encoding algorithm to operate in a batched streaming mode, enabling the processing of large-scale datasets.
@@
Batched bit-bit encoding begins by streaming the training data in batches of a specified size and incrementally fitting the dimensionality reduction algorithm. After all training batches have been processed and the dimensionality reduction model has been fit, the training data are streamed a second time. In this pass, each batch is encoded using the fitted dimensionality reduction scheme, after which batch-level mutual information importance scores are computed and the minimum and maximum values for each feature are recorded. The overall feature importance scores are obtained by averaging the importance scores across all batches. In a final streaming pass, the training data are re-encoded, min–max scaled, and written to disk along with the fitted dimensionality reduction model, feature importance scores, and per-feature minimum and maximum values. The testing data are then streamed and transformed using the saved models and statistics. To compute $Q_{\text{dataset}}$, both training and testing batches are discretized according to the specified number of qubits and the saved feature importance scores. As the training and testing batches are streamed, unique bit strings, the counts of class labels associate with each bit string, and a count of the total number of samples are stored in memory, allowing collision incidences to be computed in roughly the same manner as in the non-batched setting. The main difference is that the test-train collisions are calculated using the most common class label of the test bit string, rather than the true label of the test sample.
We study how $Q_{\text{dataset}}$ scales with both the number of features and the number of samples in a single large dataset. To this end, we subsample from the Tahoe dataset [22], which comprises approximately $100$ million transcriptomic profiles collected across $50$ cancer cell lines and $379$ drugs at varying dosages (corresponding to $1,100$ small-molecule perturbations). The Tahoe-100M dataset [22] is a unique resource created and shared through a collaboration between Vevo Therapeutics and Parse Biosciences’ GigaLab. It was intentionally generated to assist in the development of AI and machine learning tools in the domain of biology. Historically, most developers had to make do with data generated for other purposes which commonly introduces issues such as batch effects and lack of standardized metrics and annotations that accompany large biological datasets created through aggregation from different sources. Containing millions of samples across many cell lines, with each sample having measurements across thousands of interdependent features in response to different compounds delivered at varying dosages, makes this dataset ideal for exploring the applicability and scaling of quantum machine learning algorithms to understanding chaotic, non-linear, high-dimensional biological systems.
@@
In practical terms, this dataset consists of $100$ million samples, each represented by $62,710$ features (gene expression measurements). For our analysis, we restrict the dataset to a single cell line, NCI-H295 (RRID:CVCL_0456), which is the most frequently seen in the Tahoe dataset at $6\%$ of the samples. We formulate the corresponding prediction task as a multiclass classification problem: given the gene expression profile of a sample, identify which drug was applied to the NCI-H295 cell line. After filtering, the resulting Tahoe subset contains $6,040,372$ samples, $52,783$ features, and $65$ classes (the drugs applied to the NCI-H295 cell line, specifically). We then further subsample from the established Tahoe subset so we can investigate how $Q_{\text{dataset}}$ scales with the number of features and number of samples. We subsample the number of samples increasing logarithmically ($100$, $1000$, $10000$) and subsample the number of features logarithmically such that the number of features is less than the number of samples. We replicate each subsample $10$ times with a random train-test split and a biased feature selection. The feature selection is biased such that genes are chosen as features with a probability proportional to their frequency in the Tahoe subset. This is because gene expressions have been found to exhibit a power-law distribution [35], so the biased feature sampling prevent excessive sparsity in the subsample. Samples are subsampled progressively- so, for example, the subsamples with $1000$ samples will contain all the samples from the $100$ sample subsample. We obtain $Q_{\text{dataset}}$ by calculating the training and testing collision incidence when continuously increasing the number of qubits allocated to the subsample in increments of $10$ until the training and testing collision are both 0. We use batched bit-bit encoding to calculate $Q_{\text{dataset}}$, where the batch size is set to the number of samples divided by $10$.
Figure 3 shows the mean $Q_{\text{dataset}}$ over $10$ replicates as the number of features and number of samples increase. A lower number of features with the same number of samples leads to a higher $Q_{\text{dataset}}$ for the Tahoe subsamples. This trend may arise because, with many samples but relatively few features, the samples are densely packed into a smaller feature space, which increases the likelihood of collisions between samples of different classes. To resolve these collisions, the encoding requires greater bit precision per feature, leading to a higher $Q_{\text{dataset}}$. By contrast, when more features are available, samples are more spread out in the feature space, reducing collisions and therefore lowering $Q_{\text{dataset}}$. For all number of samples and number of features we tested, the mean $Q_{\text{dataset}}>50$, suggesting the learning problem on this dataset would benefit from quantum advantage. This finding reflects a growing pattern in which quantum machine learning is emerging as a potentially powerful tool for addressing biological challenges [36, 37].

### v2

@@ 4 Techniques & Results: Bit-bit Encoding Schemes
Bit-bit encoding maps numeric features and class labels into bitstrings for quantum classification. A necessary aspect of this process is compressing the original dataset. In this section, we empirically investigate how many qubits are required to represent bit-bit encoded datasets for the purpose of the learning task. We consider training and testing sets separately when calculating coverage. We calculate the training set coverage as 1training collision incidence1-$\text${training collision incidence}, where the training collision incidence is the proportion of training samples with different class labels which have the same encoding. A training set is fully covered when there are no collisions between samples with different class labels, that is, all samples with the same encoding have the same class label. In this case, since there are no colliding samples with different classes, a perfect model would theoretically be able to achieve $100\%$ training accuracy. Therefore, we refer to 1training collision incidence1-$\text${training collision incidence} as the theoretical training accuracy. We define the testing set coverage as 1testing overlap incidence1-$\text${testing overlap incidence}, where the testing overlap incidence is the proportion of testing samples which overlap with the training data but do not have the same class label as the majority class label of the overlapping training data. A testing set is fully covered when all of the samples either 1.) are not overlapping with the training data or 2.) have the same class label as the majority class label of the overlapping training data. We refer to 1testing overlap incidence1-$\text${testing overlap incidence} as the theoretical testing accuracy. We define the number of qubits at which the training and testing sets are covered (i.e. $100\%$ theoretical accuracy) as $Q_{\text{train}}(1.0)$ and $Q_{\text{test}}(1.0)$, respectively. The number of qubits required to cover the class labels is $Q_{y}=\lceil\log_{2}{(\text{number of classes})}\rceil$. Then, the number of qubits to cover an entire dataset is $Q_{\text{dataset}}(1.0)=\max{(Q_{\text{train}}(1.0),Q_{\text{test}}(1.0))}+Q_{y}$. The number of qubits required to reach an arbitrary accuracy $x$ is defined as $Q_{\text{dataset}}(x)=\max{(Q_{\text{train}}(x),Q_{\text{test}}(x))}+Q_{y}$.
While in the bit-bit encoding algorithm, the number of components after dimensionality reduction, $D$, is a modifiable parameter, unless noted otherwise, we set $D$ equal to the original number of features.
4.1 Benchmark Datasets with Principal Component Analysis
@@
Fig. 1: The train and test theoretical accuracy (left), test-train overlap (middle), and correctly classified test-train overlap (right) over the number of allocated qubits for the MNIST_784\_784 OpenML dataset with PCA dimensionality reduction. Each line shows the mean of the corresponding metric with $95\%$ confidence interval error bands. The gray dotted line is the mean $Q_{\text{dataset}}(0.99)$ (minus the class qubits). In the rightmost plot, the dip in the testing accuracy between $10$ and 2020 qubits is caused by shifts in the distribution of bits allocated to the individual features. Between $20$ and $50$ qubits, test-train overlap is near but not quite 0, so the small number of overlapping samples lead to greater shifts in the overall proportion of correctly classified samples.
We start with a simple example of calculating $Q_{\text{dataset}}(x)$ for the MNIST digit-recognition dataset. We use the MNIST_784\_784 dataset from OpenML [openml], which flattens the 2828x2828 images into $784$ features, where the features are the grayscale intensity of each pixel. The MNIST dataset contains $70,000$ samples and $10$ classes, one class for each single-digit number. Analyzing the behavior of bit-bit encoding on this dataset, as we do below, illustrates how the encoding can already indicate the achievable accuracy and generalization performance of a quantum model before it is trained.
@@
Figure 1, left, shows the mean theoretical accuracy across $10$ replicates of train/test splits. The maximum $Q_{\text{dataset}}(1.0)$ for MNIST _784\_784 is $63$, where the x-axis of the figure is cut-off, while the minimum $Q_{\text{dataset}}(1.0)$ is $42$ and the mean is $50$. The gray dotted line shows the mean $Q_{\text{dataset}}(0.99)$, $27$, minus the number of class qubits, $4$. By subtracting the class qubits, the gray line marks the average $\max{(Q_{\text{train}}(0.99),Q_{\text{test}}(0.99))}$ (recall that the theoretical train accuracy is the same as $Q_{\text{train}}(x)$, similarly for test). We postulate that large difference between $Q_{\text{dataset}}(1.0)=63$ and $Q_{\text{dataset}}(0.99)=27$ reveals a trade-off between achieving the maximum theoretical training accuracy and obtaining a generalized model. We see in the middle plot of Figure 1 that as the number of qubits increases, the overlap between the training and testing sets decreases. Between $Q_{\text{dataset}}(0.99)$ and $Q_{\text{dataset}}(1.0)$, specifically, the test-train overlap is near but not quite zero. Thus, between $Q_{\text{dataset}}(0.99)$ and $Q_{\text{dataset}}(1.0)$, the discretization of the MNIST components are becoming more and more refined for only a few samples, which reduces the overall generalizability of the encoding to new samples, or in other words, is overfitting. Furthermore, as seen in Figure 1, right, as the number of qubits increase, the test samples that do overlap with train increasingly have the same class label as the majority label of the train samples. Between $Q_{\text{dataset}}(0.99)$ and $Q_{\text{dataset}}(1.0)$, there are so few samples overlapping between test and train that slight shifts in the encoding causes the region of stochasticity in the plot.
@@
Next, we calculate $Q_{\text{dataset}}$ for a suite of benchmark machine learning datasets [feurerauto2021]. We filter the benchmark datasets to those that fulfill the following criteria: less than $1,000$ features which are all numeric, less than $10,000$ samples, more samples than features, no missing values, single-label classification with less than $10$ classes, and no collisions when not encoded. After filtering, we have $28$ datasets. Supplementary Figure S1 visualizes the number of samples, number of features, and number of class labels of each dataset. We randomly split each dataset into $80\%$ training and $20\%$ testing, repeated across $10$ replicates. We find the mean $Q_{\text{dataset}}$ for the benchmark datasets to be $27.02$ qubits, $95\%$ CI $[26.11,27.94]$. Supplementary Figure S2 shows the theoretical accuracy as the allocated bits increase for the training and testing sets. We see each dataset converges to its $Q_{\text{dataset}}(1.0)$ at a different rate and that there is minimal stochasticity in the theoretical accuracy across replicates.
We also investigate the effect of different bit-bit encoding parameters under PCA on $Q_{\text{dataset}}(1.0)$. In Supplementary Figure S3, we see that without the copula transform step, the mean $Q_{\text{dataset}}(1.0)$ of the benchmark datasets is $40$, while with the copula transform, it is $27$. The copula transform spreads samples more evenly across the discrete bins, so fewer cross-class samples fall into the same encoded bitstring and fewer test samples fall into training encodings with a conflicting majority label. In Supplementary Figure S4, we see that the mean $Q_{\text{dataset}}(1.0)$ of the benchmark datasets is largely unaffected by setting the number of components, $D$, to a percentage of the true number of features. Yet, in Supplementary Figure S5, we see whether the mean $Q_{\text{dataset}}(1.0)$ of individual datasets is higher or lower with fewer PCA components varies. On one hand, OpenML dataset “wall-robot-navigation”’s mean $Q_{\text{dataset}}(1.0)$ is $42$ bits higher with $24$ components (the number of features) than with $4$ components ($80\%$ reduction). On the other hand, OpenML dataset “steel-plates-fault”’s mean $Q_{\text{dataset}}(1.0)$ is $20$ bits lower with $33$ components (the number of features) than with $26$ component ($80\%$ reduction). We hypothesize that datasets with redundant or low-quality features exhibit lower $Q_{\text{dataset}}(1.0)$ with a smaller dimensionality, while datasets where class information is distributed across many low-variance directions will exhibit higher $Q_{\text{dataset}}(1.0)$ with a smaller PCA dimensionality.
4.2 Benchmark Datasets Across Dimensionality Reduction Schemes
@@
We now compare $Q_{\text{dataset}}(1.0)$ for the benchmark datasets [feurerauto2021] across different dimensionality reduction schemes. We include a null comparison where we don’t transform the data with a dimensionality reduction scheme (NONE). We compare three linear schemes: PCA, Independent Component Analysis (ICA), and truncated singular vector decomposition / latent semantic analysis (LSA). We also compare two non-linear schemes: radial basis function kernel PCA (KPCA) and uniform manifold approximation and projection (UMAP). All schemes, except UMAP [umap], are implemented with Scikit-Learn [scikit-learn]. Some dimensionality reduction schemes require special processing of the data — for KPCA, the training and testing sets are min-max normalized before reduction, and for ICA, the number of components is set to the minimum between the number of features and the rank of the training data.
@@
Fig. 2: Violin plots of $Q_{\text{dataset}}(1.0)$ for OpenML benchmark datasets across dimensionality reduction schemes. Each violin depicts the probability density of $Q_{\text{dataset}}(1.0)$ from the mean $Q_{\text{dataset}}(1.0)$ of each benchmark dataset.
Figure 2 shows the distributions of $Q_{\text{dataset}}$ across dimensionality reduction schemes. Overall, ICA has the lowest rounded mean $Q_{\text{dataset}}$ at $26$, followed by PCA ($27$), LSA ($28$), KPCA ($29$), NONE ($51$), and UMAP ($90$). The long tails are due to an outlier dataset in the suite of benchmark datasets — the “wall-robot-navigation” OpenML dataset [openml]. While each other dataset in the suite has a $Q_{\text{dataset}}(1.0)<50$ under PCA dimensionality reduction, “wall-robot-navigation” has a mean $Q_{\text{dataset}}=94$. This indicates significant overfitting is required to obtain full accuracy on this dataset, likely due to a few difficult-to-classify samples; for further discussion, see Section 4.3.
PCA, LSA, and KPCA transform the original feature space into uncorrelated orthogonal components, while ICA produces statistically independent components. Independent features are vital for the bit allocation step of bit-bit encoding. To see this, let $X=[x_{1},\dots,x_{D}]$ denote the $D$ components after dimensionality reduction and $y$ the class labels. Bit-bit encoding allocates bits to a component $i$ proportional to $I(X_{i};y)$, the mutual information between component $i$ and class labels $y$. If the components are not independent, they will have overlapping information, i.e. $I(X_{i};X_{j})>0$, which is over-counted in the bit allocation. The over-count can be written as $\sum_{i=1}^{D}{(I(X_{i};y))}-I(X;y)$, where $I(X;y)$ is the joint mutual information between the set of components and the class. Note that even with independent components from dimensionality reduction schemes such as PCA, there still may be unoptimal distribution of bits to components, meaning that $Q_{\text{dataset}}(x)$ is not a theoretical lower bound. The unoptimal bit distribution comes from dimensionality reduction schemes not making the components independent with regards to the class. The components may appear to be independent when not considering the class ($I(X_{i};X_{j})=0$), but in reality have synergist information with the class ($I(X_{i};X_{j};y)>0$) and redundant information with each other ($I(X_{i};X_{j}|y)<0$) or vice versa [timme2014synergy].
@@ 4.3 Results on Resource Estimation
Fig. 3: The pairwise relative differences between $Q_{\text{dataset}}(0.99)$ and $Q_{\text{dataset}}(1.0)$, $\frac{Q_{\text{dataset}}(1.0)-Q_{\text{dataset}}(0.99)}{Q_{\text{dataset}}(0.99)}$, for each benchmark dataset.
In Section 4.1 we saw a dramatic difference in $Q_{\text{dataset}}(0.99)$ and $Q_{\text{dataset}}(1.0)$ for the MNIST dataset. Figure 3 shows the relative difference between $Q_{\text{dataset}}(0.99)$ and $Q_{\text{dataset}}(1.0)$ for each benchmark dataset [feurerauto2021]. The three datasets with the highest difference are “wall-robot-navigation”, “Satellite”, and “GesturePhaseSegmentationProcessed” [openml]. In these cases, ambiguous, noisy, or contradictory samples require substantially more qubits to distinguish. A higher difference suggests that significant over-fitting would be required to train a model to $100\%$ accuracy on those datasets, highlighting a trade-off between model complexity and resource requirements, which should be considered when applying quantum machine learning as additional qubits may yield a decreased ability for a model to generalize.
@@
The universal approximation property of bit-bit encoding allows us to state that a dataset can be completely modeled with $Q_{\text{dataset}}(1.0)$ qubits. When entanglement is not bounded, classical simulation capabilities, including those that use tensor networks, plateau at around 50 qubits [quantinuum_rcs_2025]. Hence, we can claim that if $Q_{\text{dataset}}$ is $\lessapprox 50$ qubits, the corresponding dataset is unlikely to benefit from quantum advantage. Datasets that require $\gtrapprox 50$ qubits are promising targets for finding quantum advantage in machine learning within this framework. In Section 4.1, we saw that classical machine learning datasets encoded to minimize statistical dependency between features had a mean $Q_{\text{dataset}}(1.0)=27$. These classical machine learning datasets had less than $1,000$ features, less than $10,000$ samples, and less than $10$ classes (Supplementary Figure 1). This result suggests that classic, medium-sized machine learning datasets are unlikely to benefit from quantum advantage.
4.4 Large Transcriptomic Dataset
@@
Fig. 4: A heatmap visualizing how the rounded mean $Q_{\text{dataset}}(1.0)$ of subsamples of the Tahoe dataset scale with the number of features (x-axis) and number of samples (y-axis).
@@
The default bit-bit encoding algorithm requires the entire dataset to be read into memory, which becomes impractical for large datasets. Furthermore, dimensionality reduction algorithms may not scale well as the dataset size increases, leading to intractable time and memory overhead. For example, PCA dimensionality reduction scales polynomially with dataset size. Learning problems on large datasets are more likely to benefit from quantum advantage, since large datasets should have more information that can only be encoded without collision with a significant number of qubits. Therefore, we extend the bit-bit encoding algorithm to operate in a batched streaming mode, enabling the processing of large-scale datasets.
@@
Batched bit-bit encoding begins by streaming the training data in batches of a specified size and incrementally fitting the dimensionality reduction algorithm. After all training batches have been processed and the dimensionality reduction model has been fit, the training data are streamed a second time. In this pass, each batch is encoded using the fitted dimensionality reduction scheme, after which batch-level mutual information importance scores are computed and the minimum and maximum values for each feature are recorded. The overall feature importance scores are obtained by averaging the importance scores across all batches. In a third streaming pass, a bin-based copula transform model is fit on the data. In a final streaming pass, the training data are re-encoded, min–max scaled, copula transformed, and written to disk along with the fitted dimensionality reduction model, feature importance scores, copula model, and per-feature minimum and maximum values. The testing data are then streamed and transformed using the saved models and statistics. To compute $Q_{\text{dataset}}(x)$, both training and testing batches are discretized according to the specified number of qubits and the saved feature importance scores. As the training and testing batches are streamed, unique bit strings, the counts of class labels associate with each bit string, and a count of the total number of samples are stored in memory, allowing collision incidences to be computed in the same manner as in the non-batched setting.
@@
We study how $Q_{\text{dataset}}(1.0)$ scales with both the number of features and the number of samples in a single large dataset. To this end, we subsample from the Tahoe-100M dataset [tahoe], which comprises approximately $100$ million transcriptomic profiles collected across $50$ cancer cell lines and $379$ drugs at varying dosages (corresponding to $1,100$ small-molecule perturbations). Each of the $100$ million samples is represented by $62,710$ features (gene expression measurements). For our analysis, we restrict the dataset to a single cell line, NCI-H295 (RRID:CVCL_0456), which is the most frequently seen in the Tahoe dataset at $6\%$ of the samples. We formulate the corresponding prediction task as a multiclass classification problem: given the gene expression profile of a sample, identify which drug was applied to the NCI-H295 cell line. After filtering, the resulting Tahoe subset contains $6,040,372$ samples, $52,783$ features, and $65$ classes (the drugs applied to the NCI-H295 cell line, specifically). We then further subsample from the established Tahoe subset so we can investigate how $Q_{\text{dataset}}(1.0)$ scales with the number of features and number of samples. We subsample both the number of samples and the number of features logarithmically, keeping the number of features smaller than the number of samples. We replicate each subsample $3$ times with a random train-test split and a biased feature selection. The feature selection is biased such that genes are chosen as features with a probability proportional to their frequency in the Tahoe subset. This is because gene expressions have been found to exhibit a power-law distribution [gene_powerlaw], so the biased feature sampling prevent excessive sparsity in the subsample. Samples are subsampled progressively — for example, the subsamples with $1000$ samples will contain all the samples from the $100$ sample subsample. We obtain $Q_{\text{dataset}}(1.0)$ by calculating the training and testing collision incidence when continuously increasing the number of qubits allocated to the subsample in increments of $10$ until either the training and testing collision are both 0 or we reach $400$ qubits. We use batched bit-bit encoding to calculate $Q_{\text{dataset}}(1.0)$, where the batch size is set to the number of samples divided by $10$.
Figure 4 shows the mean $Q_{\text{dataset}}(1.0)$ over $3$ replicates as the number of features and number of samples increase. We see that $Q_{\text{dataset}}(1.0)$ scales with the number of samples, and below $1,000,000$ samples, $Q_{\text{dataset}}(1.0)$ stays constant with the number of features. At less than $1,000,000$ samples, the number of samples is smaller than the effective number of distinct bit strings determined by the allocated bits, so collisions are rare and $Q_{\text{dataset}}(1.0)$ is largely insensitive to the number of features. At $1,000,000$ samples, $Q_{\text{dataset}}(1.0)$ scales with the number of features; collisions begin to appear because it is more likely for multiple samples to map to the same bit strings. In this case, adding more low-information features reduces the bits allocated to informative features, increasing collisions, whereas a smaller set of highly informative features receives more bits per feature, allowing samples to be encoded more uniquely and reducing collisions. This result emphasizes the importance of feature selection to minimize the influence of low-information features in the encoding.
In Section 4.1, we saw that benchmark machine learning datasets, which had <10,000<10,000 samples and <1000<1000 features, had a mean $Q_{\text{dataset}}(1.0)=27$. We see the Tahoe subsamples at the same dimensionality have a higher mean $Q_{\text{dataset}}(1.0)=51$. The transcriptomic subsets requiring more qubits than similarly-sized benchmark datasets reflects a growing pattern in which quantum machine learning is emerging as a potentially powerful tool for addressing the unique complexity present in biological learning problems [emani2021quantum], [kubal2025quantumplatformmultiomicsdata].